OpenStudy (anonymous):

i need help with this one integrate x^x from 0 to 1

OpenStudy (anonymous):

i think i did it before for one of the users let me search for it

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^x dx\] ?

OpenStudy (anonymous):

yes thats it

OpenStudy (lgbasallote):

\[\int x^x dx\] let u = x^x du = side solution: a = x^x ln a = x ln x 1/a a' = 1 + ln x a' = x^x(1+ln x) so du = x^x (1+ ln x) dv = dx v = x \[\int x^x dx = x \cdot x^x - \int x \cdot x^x (1+ \ln x)dx\] \[\int x^x dx = x^{x+1} - \int x^{x+1}(1+\ln x)dx\] this is getting complicated

OpenStudy (anonymous):

there is no anti derivative for x^x

OpenStudy (anonymous):

here it is plz let me know if u dont get it

OpenStudy (anonymous):

no possible solution in terms of analytical form possible

OpenStudy (anonymous):

in analytical way it will get complicated can be solved numerically .

OpenStudy (anonymous):

oh...nice.... mukushala i think i should just follow your steps anyway thank u

OpenStudy (anonymous):

@moachi you know numerical methods ? like mid point rule simpson rule ? if you know i can solve it numerically.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@sami-21 yes but i wanted to know if there is any analytical solution for this or not!

OpenStudy (anonymous):

http://www.numberempire.com/definiteintegralcalculator.php and i got \[1-\frac{1}{2^2}+\frac{1}{3^3}-\frac{1}{4^4}+... \approx 0.7813\]

OpenStudy (anonymous):

using simpsons rule better accuracy can be achieved using larger n

OpenStudy (anonymous):

and thats fine with numerical computation...